OpenStudy (anonymous):

You work at a local grocery store. Your paychecks over the past 6 weeks were $56.05, $41.50, $22.50, $52.90, $49.85 and $31.00. Find the mean, median and mode. Question 1 options: mean = $45.68 median = $42.30 mode - none mean = $42.30 median = $37.70 mode - none mean = $42.30 median = $45.68 mode - none mean = $42.30 median = $45.68 mode - $31.00

OpenStudy (anonymous):

@automaticloveletter

OpenStudy (anonymous):

@tacotime

OpenStudy (anonymous):

@pooja195 (Global Intern)

pooja195 (pooja195):



OpenStudy (anonymous):

soooo????

pooja195 (pooja195):

Well let's start with the "mean" \[ \rm~\bf~56.05+41.50+ +22.50+52.90+ 49.85+ 31.00=? \]

OpenStudy (anonymous):

dead

pooja195 (pooja195):

?

OpenStudy (anonymous):

i have no idea

pooja195 (pooja195):

Well you can use a calculator .-.

OpenStudy (anonymous):

if u love me you will give me the answer

pooja195 (pooja195):

I love you as a friend but nah i'd rather have you learn. Soz.

OpenStudy (anonymous):

ok so what u want me to add

pooja195 (pooja195):

Well let's start with the "mean" \[\rm~\bf~56.05+41.50+ +22.50+52.90+ 49.85+ 31.00=?\] You can add it step by step 56.05+41.50=?

OpenStudy (anonymous):

97.55

pooja195 (pooja195):

Good continue adding the other numbers

OpenStudy (anonymous):

75.40

pooja195 (pooja195):

Ok when you add up the numbers you get 253.8 Now we need to divide by the number we added. How many numbers did we add?

OpenStudy (anonymous):

idk yungin

pooja195 (pooja195):

-.- if you don't start putting in some work i'll leave

OpenStudy (anonymous):

ill put dat work in you

pooja195 (pooja195):

Bye -_-

OpenStudy (anonymous):

you just wont get a medal

pooja195 (pooja195):

I dont care about medals :P now if you'd really like help do the work.

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

what do i divide

pooja195 (pooja195):

Ok when you add up the numbers you get 253.8 Now we need to divide by the number we added. How many numbers did we add?

OpenStudy (anonymous):

4

pooja195 (pooja195):

count again.

OpenStudy (anonymous):

its a. right

OpenStudy (anonymous):

i gotta hurry this question up no disrespect

OpenStudy (anonymous):

anybody knw the answer

pooja195 (pooja195):

It's not A.

OpenStudy (anonymous):

b.

pooja195 (pooja195):

-_- im done bye.

OpenStudy (anonymous):

you would just leave me

pooja195 (pooja195):

;)

OpenStudy (anonymous):

whats all the numbers i got to divide

pooja195 (pooja195):

56.05, 41.50,22.50,52.90, 49.85,31.00 We have 6 numbers we added so we are dividing \[\huge~\rm~\bf~\frac{ 253.8 }{ 6 }=? \]

OpenStudy (anonymous):

ok bet

OpenStudy (anonymous):

42.6

pooja195 (pooja195):

Wrong

pooja195 (pooja195):

the decimal is wrong

OpenStudy (anonymous):

3

pooja195 (pooja195):

42.3 = mean so: Mean=$42.30

pooja195 (pooja195):



OpenStudy (anonymous):

so it is b. right

pooja195 (pooja195):

Now to find the mode \[\rm~\bf~56.05, 41.50,22.50,52.90, 49.85,31.00\] @milliaunjiral quit guessing >_>

OpenStudy (anonymous):

ight

OpenStudy (anonymous):

there is none

pooja195 (pooja195):

Well done ok next find the median The median of a data set can be found by arranging all the values from lowest to highest value and picking the one in the middle.